OpenStudy (anonymous):

What is the slope of the line given by the following equation? y = -4x + 9

hero (hero):

y = mx + b m = slope y = -4x + 9 -4 = slope

OpenStudy (anonymous):

wow thank you

hero (hero):

In general, if y = mx + b, m is the slope In this particular case, if y = -4x + 9, then -4 is the slope.

hero (hero):

That's about the simplest way it can be put